HI and H_{2} in Luminous Interacting Galaxies

N. Scoville^{1}, J.E. Hibbard^{2}, M.S. Yun^{1}, J.H. van Gorkom^{2}

1California Institute of Technology, Astronomy Department 2Columbia University, Astronomy Department

ABSTRACT

The interstellar gas plays a fundamental role in the interacting galaxy systems: it is dissipative and thus responds irreversibly to strong dynamical perturbations; it is also the active component which fuels starbursts and may fuel AGN’s in the central regions of the galaxies. In this article we review new high resolution aperture synthesis observations of the atomic and molecular gas in luminous interacting galaxies.

Due to the extended nature of the atomic gas in disk galaxies, the HI observations bear most critically on the long range tidal perburbations and therefore serve to elucidate the interaction history. Here we summarize the HI observations of six systems: the M81/M82 group and five interacting pairs with long tidal tails. In all of these systems the impact of geometry is evident from the spatial and kinematic structure of the atomic gas within the extended tidal features. The six systems may be crudely placed in a temporal sequence for galactic merging – in the early stages large amounts of HI still exist within the galactic disks and the star formation is spread throughout the disks as evidenced by Hα emission while in the final stages the HI is almost entirely contained within the tidal features beyond the merger body and the molecular gas is mostly found in the central remnant.

Molecular line observations have largely concentrated on the intermediate evolutionary phase in which the galaxies become strongly emitting in the far infrared. Molecular line surveys of luminous infrared bright galaxies sampled by the IRAS survey have shown these galaxies to be extraordinarily rich in molecular gas as evidenced by the CO emission. High resolution aperture synthesis data are presented for two of these systems – also representing a temporal sequence in the intermediate phases of merger induced activity. In each case the molecular gas is found to be much more highly concentrated than in normal galaxies, probably a result of the viscous dissipation of energy and angular momentum in the ISM as a result of the strong dynamical perturbations. Within the central few hundred parsecs as much as 10^{10}M_{⊙} of H_{2} is found. Such concentrations are surely the source of nuclear starbursts, possibly also rejuvinating nuclear nonthermal activity. It is suggested that interaction induced activity in the interstellar gas of galaxies may play a fundamental role in galactic evolution – leading to star formation rates elevated by at least an order of magnitude for a period ∼ 2 × 10^{8} years. Since such activity is presently occurring in approximately 2% of all spiral galaxies, it is clear that galactic interactions can be critical to both the long term stellar populations of galaxies and to the most luminous, energetic phases for galactic evolution.

1 INTRODUCTION

The important role of interstellar gas in galactic interactions has been recognized only in the last few years. It has long been suggested that interaction induced gas inflow may play an important role in the enhanced central activity (see discussions by Kennicutt and Heckman in this volume), but the dynamical importance of interstellar gas in galactic interactions has been recognized only recently. The critical difference between the gas and the stars is that the gas clouds have a relatively short mean free path and are highly dissipative. The gas is also the medium from which luminous young stars form. The atomic hydrogen in the extended tidal features provides us with a unique record of the galactic encounter from its spatial distribution and kinematics.

The last few years have seen an explosive growth in high resolution observations of both the low density atomic and high density molecular (via CO) phases of the ISM in interacting galaxy systems. Much of the interest in these systems was spurred by follow-up studies from the IRAS survey revealing a strong correlation between infrared luminous galaxies (L_{IR} > 10^{10}L_{⊙}), presumably undergoing bursts of star formation, and optical morphologies suggesting a past or ongoing galactic interaction. Under the assumption that most of the luminosity seen in the far infrared originates as UV and optical radiation from young stars which has been absorbed and reradiated by dust, the far infrared luminosity provides an indication of the present day star formation rates, albeit with uncertainties related to the form of the stellar inital mass function. Recent high resolution near infrared imaging has pinpointed the centers of activity — most often in the nuclei of the merging galaxies. Relative to noninteracting galaxies of similar interstellar gas content, the inferred star formation rates are elevated by an order of magnitude and occasionally by two orders of magnitude.

The importance of understanding interacting galaxy systems arises from the dramatic effect these interactions may play in determining large scale galactic structure and in resuscitating a most luminous phase in galactic evolution. Mergers were probably more frequent in the past and they might be responsible for much of the present structure of galaxies. In this sense the study of galactic mergers can shed light on protogalactic evolution and hierarchical formation of galaxies (e.g. White & Rees 1978, Peebles 1980).

In this review we present HI and H_{2} data for a sample of interacting systems which range in sequence of elevated activity and probably also temporal evolution of galactic merging. The HI data often reveal extended tidal features not at all apparent in optical images while the CO observations reveal massive concentrations of gas which are presumably the sites of interaction-induced starbursts. As the interaction progresses the atomic gas resides predominantly in the tidal features outside the remnant body, while the molecular gas appears concentrated in the nuclei of the merging systems. Using canonical conversion factors from CO to H_{2}, the molecular gas often accounts for more than 30% of the indicated dynamical mass in the central regions. Since this is much higher than the mass fractions found in the presumed progenitors (< 10%), it is a strong indication that gaseous dissipation has played a role, and perhaps a dominant one, during the merger (Kormendy & Sanders 1992).

2 HI - THE INTERACTION RECORD

Neutral hydrogen studies are particularly useful for disentangling the geometry of interactions involving disk galaxies due to two facts: (1) the tidal features are drawn primarily from the outer regions of the disks (c.f. Toomre & Toomre 1972) which are rich in atomic gas (Wevers et al. 1986); and (2) this outer gas has not undergone more than 1-2 rotations since the encounter started and therefore retains spatial and kinematic signatures of the encounter (Toomre & Toomre 1972).

In this section we describe the observational results for six systems: the M81/M82 group, Arp 295, NGC 4676 (the “Mice”), NGC 520/UGC 957, NGC 3921, and NGC 7252 (the “Atoms for Peace” galaxy). For the M81/M82 system, a numerical model has been constructed for the encounter. For the other systems, detailed models have not yet been generated, but the signature that the interaction has left on the outer gas is clearly seen.

2.1 The M81/M82/NGC 3077 Group

At a distance of only 3.3 Mpc the M81/M82/NGC 3077 group is the closest of those studied at high resolution. It differs from those discussed later in that it is dominated by a single massive galaxy, M81, with two smaller companions. The entire group has recently been mapped with the VLA D array at 60′′ resolution by Yun, Ho & Lo (1993a), and the mosaic HI image of twelve fields is shown in Figure 1. For comparison at the same scale an optical photographic image is reproduced in Figure 1b. The comparison most clearly demonstrates the unique ability of the HI emission to trace extended tidal features well beyond the optical disks. The total atomic hydrogen mass in the group is approximatley 6 × 10^{9}M_{⊙} with 60% contained in the disk of M81 and 20% concentrated near NGC 3077 to the southeast. The remainder of the HI, approximately 10^{9}M_{⊙}, is contained in the extended spiral/tidal arm reaching from M81 through NGC 3077 and up towards M82 to the north.

Yun (1992) modeled these data by tracing the response of the disk particles under the influence of the potentials representing three orbiting galaxies. As shown in Figure 2, the hyperbolic passage of M82 from the south and the parabolic passage of NGC 3077 from the northeast result in tidally driven spiral arms in M81 and severe disruption of the disks of M82 and NGC 3077. Although most of the HI seen outside the disks of the galaxies probably originated from M81, one exception is the apparent tidal tail extending from NGC 3077 to the northwest, tracing the orbit of M82. This feature would be hard to understand unless the gas has originated from NGC 3077, and thus all three galaxies probably had substantial amounts of gas prior to the interaction.

The major effects of the interactions among the three galaxies in this group are: the stimulation of the strong spiral density wave in the disk of M81; a complete disruption of the disk of NGC 3077; and the central concentration of gas in the disk Figure 1a A VLA D array map of the integrated HI emission in the M81 group is shown in grey scale. In the region mapped, the total HI mass is 6 × 10^{9} M_{⊙}, 60% of which is contained in the disk of M81.

of M82. With regard to the strong spiral pattern in M81, it is interesting to note that the presence of four spiral arms is probably consistent with there having been two interactions (i.e. from both NGC 3077 and M82). The numerical simulation required a nearly simultaneous interaction involving all three galaxies to reproduce the tidal features found around NGC 3077.

The higher resolution HI observations of M82 using the VLA C array (Yun, Ho & Lo 1993b) found an amorphous disk and a pair of prominent tidal tails extending out of the stellar disk, indicating a severe tidal disruption of its own HI disk rather than accreting gas captured from M81. The HI kinematics suggests the presence of a bar potential in M82, which may be responsible for channeling the gas into the nuclear region. The extraordinary effects on M82 are best seen in molecular line Figure 1b Optical photographic image of the same area in the M81 group shown in the Figure 1a HI image.

tracers and in the infrared. In this galaxy, the total molecular gas mass may be as much as 25% of the dynamical mass and the dense gas is highly concentrated in the center (cf. Young & Scoville 1984). The nucleus of M82 has been mapped at high resolution in CO, HCN and HCO^{+} by Lo et al. (1987), Nakai et al. (1987), and Carlstrom (1990). These studies reveal a torus of gas at radius ∼ 150 pc and mass ∼ 10^{8}M_{⊙}. The relatively strong emission from high dipole moment molecules such as HCO^{+} and HCN (Carlstrom 1990) suggests that the physical conditions in this gas are quite different from those in typical Galactic giant molecular clouds (GMCs). The derived kinetic temperatures are 50K and the volume densities within the clouds are approximately 10^{5} cm^{−3}. The fact that the bulk of the near infrared stellar light, presumably from relatively young giant stars, occurs within the torus has led to the suggestion that the starburst is progressing from the center outwards. Carlstrom Figure 2 A schematic illustration of the behavior of disk particles in M81, M82, and NGC 3077 from the numerical simulation by Yun (1992). The orbits of the galaxies are traced in solid lines. The second, third and fourth panels show separately the particles from each individual galaxy in its rest frame.

(1990) estimates from the 3 mm free-free emission that the UV production rate is 6 × 10^{53} sec^{−1}, corresponding to 3 × 10^{5} O8 stars. This ionizing photon production rate deduced for the center of M82 (R < 500 pc) is in fact somewhat greater than that of the entire Milky Way. A recent analysis of the cosmic ray heating for the moleuclar gas in M82 by Suchkov, Allen & Heckman (1993) suggests that elevated temperatures seen in the molecular gas can be understood if the cosmic ray density is approximately 300 times that of the Milky Way as would be expected in such a starburst.

2.2 HI in Strongly Interacting Systems

The systems studied at high resolution with the VLA by Hibbard et al. (1993a) and Hibbard & van Gorkom (in preparation) were chosen for the appearance of their extended tidal tails. Four of the five galaxies were taken from Toomre’s (1977)

Figure 3 The integrated HI intensity from the VLA C+D observations of Arp 295 is shown superposed as contours on a negative grayscale of the KPNO 2.1m R-band image with all pixels containing Hα emission set to white.

proposed sequence of approximately equal mass disk galaxies in progressive stages of merging. The goal of the study is to quantify the atomic gas distribution and star formation activity as one progresses in a sequence through more advanced interactions and merging.

Though they were selected purely on the basis of their optical appearance, all of these systems turned out to be IR luminous, with IR luminosities in the range of 10^{10} − 10^{11}L_{⊙} (H_{0} = 75 km s^{−1} Mpc^{−1} is used throughout), and all show signs of central activity: all are radio continuum sources, and all but Arp 295B and NGC 7252 have optical spectra typical of LINERs (Keel et al. 1985, Dahari 1985.

Added to the beginning of the Toomre (1977) sequence is Arp 295, at a distance Figure 4 HI intensity contours from the VLA C+D array observations are shown superposed on a KPNO 2.1m optical (negative greyscale) and Hα (white) images of the NGC 4676 system, “The Mice”. The inset shows a close-up of the disks of both galaxies, with the optical image in negative greyscales and the Hα emission shown as contours upon a white background. The plume of Hα emission along the minor axis of the northern galaxy and the offset optical and gaseous bar of the southern galaxy are seen in this inset.

of 88 Mpc. In Figure 3 the integrated HI intensity is shown contoured on a negative greyscale of the R-band optical light with white pixels indicating Hα emission. The system has both a bridge and a tail indicating at least one pass of the two galaxies about each other, but both disks are still distinct indicating that the interaction is not too far advanced. The HI observations reveal that this system is part of a loose group, with nine smaller galaxies detected in HI, three of which are seen in Figure 3.

Already at this relatively early stage of interaction we find approximately 30% of the total HI (M_{tot}(HI) = 3×10^{10} M_{⊙}) outside of the optical bodies of the galaxies. The bridge extending approximately 120 kpc between the two galaxies is continuous in the optical continuum from stars but is broken in HI. The Hα emission is widely distributed through both disks and there is no evidence of extensive star formation occurring within the tidal features. Plumes of Hα emission are also seen along the minor axis of the northern disk galaxy.

The next system in the sequence, NGC 4676 (the “Mice”), is at approximately the same distance as Arp 295. The optical continuum, Hα, and neutral hydrogen emission are shown in Figure 4. In Figure 4a, extended tidal tails are seen on both sides of the two galaxies and a luminous bridge of material flows between them. Approximately 65% of the total HI mass (9×10^{9} M_{⊙}) appears outside of the optical disks. The Hα is widely distributed over both galactic disks and tidal tails and along the minor axis of the northern companion. The details of the disks are seen in Figure 4b, with Hα shown contoured on a white background, and the optical light in negative greyscales. This figure shows very clearly the Hα along the minor axis of northern galaxy, as well as a bar in the southern galaxy at approximatly PA∼ 30 deg, both in the gas and the stellar light. The gaseous bar leads the stellar bar as seen in the numerical calculations of Barnes & Hernquist (1991) who point out that the offset stellar bar will torque the gas, leading it to loose its angular momentum.

The system consisting of NGC 520 and UGC 957 is at a distance of approximately 30 Mpc. This system is a more advanced merger than the previous systems, and a near infrared (2μm) map clearly shows the presence of two nuclei (Stanford & Balcells 1991) — a primary one coinciding with the center of the dust lane, and a secondary one to the northwest, still within the main body of the NGC 520. At the location of the primary nucleus is a strong extended radio-continuum source (Condon et al. 1982) and a molecular gas disk (Sanders et al. 1988a). This gas concentration has a column density that is ten times that of typical Galactic GMCs, and comprises 24% of the indicated dynamical mass. The Hα emission is distributed throughout the main body of NGC 520: at both nuclei, along the minor axis of the primary nucleus, and within the inner tidal loops and tails.

UGC 957 is seen much further to the northwest (∼ 55 kpc) within the center of a long HI arm/tail, much like NGC 3077 in the M81 system (Figure 1). But unlike in M81, the kinematics of the outermost HI in NGC 520 is dominated by rotational motion, rather than swinging tidal motions (c.f. van Gorkom 1993). At present 60% of the HI (9×10^{9} M_{⊙}) is outside the main optical disk, and much of this gas has no optical counterpart.

NGC 3921 (=Mkn 430) consists of a single nucleus with two tidal systems — one stellar and one gaseous — suggesting that this system may be the result of a merger of a spiral and a gas poor S0 galaxy. That the stellar light distribution is well described by an r^{\frac{1}{4}} law (Stanford & Bushouse 1991) is an indication that the system is in a late stage of merging and has relaxed in the inner regions. This is supported by spectroscopy of the stellar light from the nucleus, which is consistent with a starburst approximately 10^{9} years ago (Schweizer 1990). The presence of a large star-forming cluster in the southern tail suggests that self gravitating systems may well be formed in some of the tidal tails of interacting galaxies (c.f. Hibbard & van Gorkom 1993). In NGC 3921, 75% of the HI (7×10^{9} M_{⊙}) is located within the tidal features, and the kinematics suggests that perhaps all of it is outside of the main body of the remnant. The Hα emission in the body is more regular than in the earlier stages, and extensive ongoing star formation is seen along the southern tail even at this late stage.

NGC 7252 (the “Atoms-for-Peace” galaxy) is probably an advanced merger of two disk systems. The optical light distribution indicates an r^{\frac{1}{4}} law, and the merged galaxy may be well along towards forming an eliptical galaxy core (Schweizer 1982). This conclusion is supported by the ageing burst population evident in the optical spectra (Schweizer 1982). NGC 7252 has also been mapped in CO by Wang, Schweizer, & Scoville 1992, who find that most of the molecular gas (80% of the total CO emission) is in the very central regions. This is in contrast to the 5×10^{9} M_{⊙} of atomic gas, all of which is contained in the tidal features (Hibbard et al. 1993b). Hα emission is seen not only in the nucleus but also filling in the gap between the CO and the innermost HI. Recent ROSAT observations reveal X-ray emission from NGC 7252, possibly indicating the presence of hot gas (Hibbard et al. 1993b).

For this system, detailed observations suggest that the atomic gas appears only in the tidal tails for two reasons: the outer HI is sent to these regions by the tidal forces experienced during the merger, and any inner HI is converted to other forms. The central molecular-gas coincides with the inner star-forming disk and accounts for 45% of the indicated dynamical mass, suggesting that dissipation has played a dominant role during the merger (Kormendy & Sanders 1992). The only ongoing star formation in the tails coincides with large concentrations of atomic gas and optical light, possibly being assembled out of the tidal material, and the blue colors seen throughout most of the tails indicate a young stellar population.

Briefly summarizing the qualitative changes as one goes from the less advanced to the more advanced interacting systems: 1) a larger fraction the HI appears outside the optical bodies in the extended tidal features and the most advanced mergers show virtually no HI left within the optical bodies; 2) the Hα is at first widely distributed through the galactic disks, and frequently along the minor axis of the galaxies. At intermediate stages there are signs of increased activity in the central regions, and star formation is seen more widely in the tidal tails. Finally, in the latest stages, the star formation is centrally concentrated towards the merger nucleus, and the tails have blue colors; 3) in the later stages, large concentrations of molecular gas are collected in the remnant body, which has relaxed into a r^{\frac{1}{4}} law in its optical light and exhibits a post-starburst spectrum.

Apparently large quantities of both gas and stars are sent to large radii during such interactions, and large quantities of gas are brought into the central regions where it forms stars. The form that this material takes during its sojourn into intergalactic space and where it ends up after a few dynamical times will have important reprecussions for the future evolution of the remnants. We find star formation taking place abundantly in the tails, and bound subsystems may be forming from the tidal debris as originally suggested by Zwicky (1956). The peak column densities of atomic hydrogen within the tidal tails averaged over the resolution are at the threshold column density at which molecular hydrogen usually dominates atomic hydrogen within the disks of normal spiral galaxies, and systematic searches for molecular emission within the tidal tails might prove fruitful.

3 MOLECULAR GAS - THE ACTIVE PHASE

In recent years a major direction of extragalactic molecular line research has been in studies of infrared bright galaxies highlighted as a result of the IRAS Survey. At the same time major ground based follow-up studies have been done in the optical, near infrared, and radio continuum (e.g. Lawrence et al. 1986). The IRAS bright galaxy survey (Soifer et al. 1987) included 238 objects in the northern hemisphere. Optical imaging of this sample reveals a strong correlation between the infrared luminosities of the galaxies and the existence of morphological features suggestive of a strong galactic interaction or merger (e.g. extended tidal tails or double nuclei; Sanders et al. 1990). At the lowest luminosities, 10^{10} − 10^{11}L_{⊙}, the large majority of the IRAS galaxies appear isolated and undisturbed whereas at 10^{11}−10^{12}L_{⊙} approximately 30% show evidence of interactions and above 10^{12}L_{⊙}, ≥ 90% appear to have undergone a recent strong interaction.

The ultimate source of the IR luminosity is likely trapped UV radiation from either a central starburst or AGN, which heats the dust that has been widely distributed during the encounter. Using VLBI radio observations, Norris et al. (1990) concluded that when an AGN was responsible for the central activity, it generally made itself known through other means, such as the optical emission lines (c.f. Sanders et al. 1988b). Thus, some IR galaxies are probably powered by an AGN, but while others, especially the less luminous ones, by starbursts.

Molecular line observations of the most luminous IRAS galaxies are relevant for understanding of the effects of interactions on the molecular gas phase. In order to access the total molecular gas content of the most luminous IRAS galaxies, Sanders, Scoville & Soifer (1992) obtained single dish measurements for a sample of 89 galaxies from the IRAS bright galaxy survey. The resulting H_{2} masses, assuming the same CO emissivity per unit mass as for the Milky Way GMCs, are shown in Figure 5, plotted as a function of infrared luminosity. Also shown are the equivalent masses and luminosities of “normal” galaxies – the Milky Way and M51.

It is clear from Figure 5 that the most luminous IRAS galaxies are extremely rich in molecular gas with total H_{2} masses in the range 2 − 50 × 10^{9}M_{⊙}, that is 1–20 times the H_{2} content of our Galaxy. It is also clear from Figure 5 that these galaxies are not simply scaled up versions of normal galaxies since the infrared luminosity to H_{2} mass ratio is usually elevated considerably from the value of ∼ 4L_{⊙}/M_{⊙} seen in normal galaxies.

On the basis of their optical morphologies the galaxies in the CO survey sample were placed in one of four possible categories: isolated spirals, spirals with a disturbed appearance, galactic pairs with non-overlapping disks, and advanced mergers. The latter category was distinguished from single isolated galaxies on the basis of extended Figure 5 The infrared luminosity (λ = 8 − 1000μ m) is shown as a function of the global molecular gas mass for a sample of 89 bright infrared galaxies (Sanders, Scoville & Soifer 1990; Tinney et al. 1990). Comparable data are also shown for the Milky Way and M51.

tidal tails in the optical or double nuclei seen in the near infrared (e.g. Carico et al. 1992). In Table 1 the mean infrared luminosity-to-molecular mass ratios are given as a function of infrared luminosity and interaction classification. The number of galaxies within each category is shown in parentheses. For the purpose of the present discussion the most important result is summarized in the last column. Averaging over IR luminosity it can be seen that both the galactic pairs and advanced merger categories have approximately an order of magnitude higher IR luminosity-to-gas mass ratio than the normal galaxies. It is not totally clear from this sample that the overall level of energetic activity is ten times higher in strongly interacting systems than in isolated galaxies since this sample is biased towards bright infrared galaxies. It will be important for a future study to investigate the IR luminosity-to-gas mass ratio in a sample which is not biased towards infrared bright galaxies.

A more surprising result of the molecular line observations in the interacting

TABLE 1. Luminosity-to-H_{2} Mass Ratios

L_{IR} (10^{11}L_{⊙})
L_{IR}/M_{H_{2}} (#Gal) 0.2 − 1 1 − 10 10 − 30 all

Morphology

Isolated  6 (15)  –  –  6 (15)
Disturbed  12 (13) 22 (11)  –  17 (24)
Non-overlapping Disks  20 (4)  30 (7)  75 (1)  30 (12)
} 39(47)
Merger Pairs  32 (5)  34 (23)  74 (7)  42 (35)

L_{IR} is the infrared luminosity at λ = 8−1000 μm and M_{H_{2}} is the H_{2} mass assuming a Galactic CO-to-H_{2} conversion ratio of 3 × 10^{20} cm^{−2} (K km s^{−1})^{−1}. The units are L_{⊙}M_{⊙}^{−1} and the number of objects in each group sampled by Sanders, Scoville, and Soifer (1990) and Tinney et al. (1990) is given in parenthesis.

galaxies highlighted as a result of the IRAS survey has been the finding that the molecular gas is not only more abundant, but also more concentrated than in normal galactic disks. To date, approximatley 30 of the high luminosity galaxies have been imaged with millimeter interferometers in the CO line. In Table 2, observational results are summarized for 19 systems observed with the Owens Valley array. The galaxies listed in Table 2 span the luminosity range 2 × 10^{10} − 3 × 10^{12}L_{⊙}(λ = 8 − 100μm). Ten galaxies, some of which are in the Owens Valley sample, have also been imaged using the Nobeyama array (Okumura et al. 1991). The significant results born out in both studies are: a significant fraction of the total molecular gas, usually greater than 50%, is typically concentrated in a nuclear source less than or equal to 1 kpc in radius; and secondly, the gas mass fractions in the center (M_{H_{2}}/M_{dyn}) often exceed 50% (as compared with typical values of 5% near the centers of normal galaxies). Below, we describe in detail two cases — NGC 4038/39 and Arp 220 — which serve to illustrate molecular gas morphology in intermediate and high luminosity interacting galaxies.

3.1 NGC 4038/39 — The Antennae

At a distance of 21 Mpc, the “antennae” is one of the best examples of a strong interaction in progress where excellent spatial resolution can be achieved. Stanford et al. (1990) imaged two 1′ fields in CO at 6′′ resolution. The integrated intensity maps are shown in Figure 6, superposed on a 6500 Å image of the inner disks. Three major CO concentrations are evident — two centered near the galaxy nuclei and a third northeast of NGC 4039, where the two galaxies overlap. To within the position uncertainties, the two nuclear CO peaks coincide with 2μm peaks. The third CO complex, in the overlap region, coincides with a number of radio continuum, Hα, and 10μm emission knots (Hummel & van der Hulst 1986; Rubin, Ford, & D’Odorcio 1970; Wright et al. 1988). The CO concentration at the nucleus of the northern galaxy, Figure 6 CO emission contours are shown superposed on a 6500 Å image of the NGC 4038/39 system. In the upper right a deep, larger scale image showing the tidal tails is also shown. Three massive concentrations of molecular gas are seen: at the centers of the two galaxies and at a possible interaction interface where the two galaxies overlap. (Stanford et al. 1990).

NGC 4038, is approximately 1 kpc in diameter and the integrated CO intensity suggests a molecular mass of approximately 8 × 10^{8}M_{⊙}. At the NGC 4039 nucleus, an H_{2} mass of 2 × 10^{8}M_{⊙} is found.

The massive gas concentration in the overlap region is probably the result of cloud-cloud collisions within the disks of the two colliding galaxies. Such collisions will be highly inelastic due to rapid cooling in the dense gas, and cancellation of angular momentum may result in the eventual accumulation of much of this gas near one of the galactic centers or in a merged nucleus. The N-body simulation of the interaction by Barnes (1988) suggests a time since pericenter passage of about 2×10^{8} years, with the two disks brushing their edges at that time and the two nuclei separated by 20 kpc. The resulting gravitational perturbation could concentrate the originally more tenuous disk gas into the giant off-nuclear CO complexes which are now seen. Indeed the long northern tail pulled away from the outer disk of NGC 4039 appears to lead to the inner disk of NGC 4038, passing through the CO super association. Thus, gas in the disks might have been shocked, perhaps initiating the off-nuclear starburst.

3.2 Arp 220 – An Advanced Merger

Arp 220 has an infrared luminosity at λ = 8 − 1000μm of 1.5 × 10^{12}L_{⊙}, exceeding that in the visual by two orders of magnitude and placing it in the luminosity regime of quasars. At optical wavelengths, the galaxy appears approximately spherical with a central dust lane and faint tidal tails extending up to 70 kpc away (Sanders et al. 1988b). Single dish measurements show the CO emission spanning 900 kms^{−1} and the derived H_{2} mass is 3.5 × 10^{10}M_{⊙}, approximately a factor of 15 greater than that of the Galaxy (Solomon, Radford, & Downes 1990).

In Figure 7, maps at 2′′ resolution (750 kpc) are shown from Scoville et al. (1991). The 3 mm continuum (mostly dust emission) and CO line emission are shown in the upper panels as a function of displacement coordinates from the cm-wave radio continuum peak (Norris 1988). Both the CO and 3 mm continuum peak are within 0.5′′ of the western component of the nucleus seen in high resolution near infrared and radio maps (e.g. Graham et al. 1990). In addition to the compact nuclear source, an extended CO emission component can be seen in maps with smaller velocity ranges. In the lower panels of Figure 7, the emission is shown for velocity windows centered at 5330 and 5642 km s^{−1}. The low level contours of CO emission are clearly extended along a NE-SW direction, parallel to the dust lane.

The CO emission can be separated into two components: a 1.4′′ × 1.9′′ core and 7′′ × 15′′ extended component. The H_{2} masses in the core and extended components are 0.9 and 1.8 × 10^{10}M_{⊙} respectively. The mean diameter of the core component (∼ 1.7′′) corresponds to a radius of 315 pc and the mean H_{2} density over the volume is 2900 cm^{−3}. The dynamical mass for the core is similar to the H_{2} gas mass there.

In Figure 7 (lower panels), a velocity gradient in both the core and extended components may be seen - lower velocity emission displaced to the SW, the higher velocity emission to the NE. The direction of this velocity gradient is parallel to the major axis (PA= 52 − 63 deg) of the two CO emission components and the dust lane seen in the optical. Thus, the molecular gas has partially relaxed to a disk-like configuration with angular momentum playing a significant role in the gas distribution, probably inhibiting further collapse towards the center. The upper limit to the emission measure of ionized gas derived from the 3 mm continuum permits only 10^{11}L_{⊙} for a population of only O5 stars. If a starburst is to account for the total luminosity, its duration must exceed 10^{9} yr or the IMF must be deficient at Figure 7 The λ = 2.7 mm dust emission and integrated CO emission are shown in the upper panels as a function of offsets from the centimeter wave non-thermal radio peak in Arp 220 (Scoville et al. 1990). In the lower panels the CO emission is shown separately for a blueshifted and redshifted veloctiy ranges. The synthesized beam is 1.9 × 2.1′′.

≥ 15M_{⊙} (Scoville and Soifer 1991).

3.3 Later Evolutionary Phases

There are a number of objects mapped at high resolution in molecular lines which have not recently undergone an interaction, but which may be reasonable analogs of the late evolutionary phases for interacting galaxy systems – that is an ageing nuclear starburst or a dust obscured AGN.

NGC 3079 is an edge-on spiral galaxy exhibiting radio lobes emerging from its nucleus and a high concentration of molecular gas within the central 1 kpc. The galaxy has recently been studied at high resolution in CO (Sofue & Irwin 1992). Within the central 200 pc a rapidly rotating core of gas can be seen with total mass approximately 1.5 × 10^{9}M_{⊙} and a gas mass fraction of approximately 50%. At slightly larger radii they see approximately 10^{10}M_{⊙} of molecular gas situated in a ring feature. Sofue and Irwin (1992) also found molecular gas along the periphery of one of the minor axis radio lobes and they suggest that the nuclear radio jet is possibly focused by molecular gas. Tilanus & Veilleux (1993) compare the distribution of CO and HCN emission in the center of NGC 3079 — the latter having a higher dipole moment preferentially arises from molecular gas at density ∼ 100 times greater than that sampled in CO. They find a strong gradient in the ratio of HCN to CO, with the latter more centrally concentrated in the core. Thus, it appears that the volume densities within molecular clouds are considerably higher at small radii, i.e. similar to the situation found in M82 (§2.1).

In NGC 1068 (Seyfert II), the molecular gas is mainly distributed in two massive spiral arms which originate from the ends of a stellar bar seen at 2μm. Between these spiral arms at 1.4 kpc radius and the nucleus, there exists relatively little molecular gas, but on the nucleus there is possibly 6×10^{7}M_{⊙} of H_{2} (Planesas, Scoville & Meyers 1991). This gas (and associated dust) may hide the hypothesized Seyfert I nuclei (see Antonucci & Miller 1985). It is interesting to note that the molecular gas and young stars in NGC 3310 (Kikumoto et al. 1992) exhibit similar distributions to those in NGC 1068.

A number of observational and theoretical studies (eg. Sanders et al. 1988b, Norman & Scoville 1988) have suggested that in some cases strong galactic interactions in gas-rich galaxies might ultimately lead to the formation and fueling of a central AGN or quasar. On the observational side, the optical emission lines seen in approximately 50% of the ultraluminous infrared galaxies have intensity ratios similar to those seen in LINER and Seyfert I nuclei. In fact, strong CO emission has now been seen in four UV-excess quasars: MRK 1014, IZw 1, PG 0838+77, and Mrk 876 (Sanders, Scoville & Soifer 1988; Barvainis, Alloin & Antonnucci 1989; Alloin et al. 1992). The Owens Valley array has been used recently for a sensitive search for CO emission in the radio loud quasar 3C 48 (Scoville et al. 1993). A probable emission feature was detected at the redshift of the narrow optical emission lines (z = 0.3695) and the width of this feature is approximately 300 km s^{−1}. The integrated CO line flux implies a H_{2} mass of 7 × 10^{10}M_{⊙}. 3C 48 is a particularly significant object inasmuch as it was the first optically identified quasar (Matthews and Sandage 1963) and yet it has many other characteristics similar to those of the ultraluminous infrared galaxies. It has a strong far infrared excess (Neugebauer, Soifer & Miley 1985), extended emission line gas and a possible double nucleus (cf. Stockton & Ridgway 1991). The detection of strong Balmer absorption lines in the extended nebulosity also indicates a luminous young stellar population (Boroson & Oke 1984), possibly an ageing starburst.

Probably the most extreme object with respect to its interstellar medium mass is the recently detected object IRAS 10214+4724 at z = 2.3 (cf. Brown & Vanden Bout 1991). Here the mass of molecular hydrogen exceeds 10^{11}M_{⊙} and the far infrared luminosity is approximately 3 × 10^{14}L_{⊙}. Recent near infrared images obtained by Matthews et al. (1993) with the Keck telescope shows several near infrared peaks, possibly indicating a compact cluster of interacting galaxies.

4 GALACTIC EVOLUTION RESULTING FROM INTERACTIONS

In the previous sections we have attempted to show the important role played by the interstellar medium during galactic interactions — both in terms of the record of the encounter left in the extended tidal tails seen in HI and the dissipative/starburst role played by the molecular gas. We now address the issue of the long term effects which interactions of gas-rich galaxies may have on galactic evolution. We have seen from the molecular line data that an enormous mass of molecular gas may be concentrated at the center of the interacting system, while the diffuse atomic gas is relagated to the extended tidal tails. In the nucleus the star formation will be both more concentrated and much more rapid as judged from the infrared spatial distribution and the IR luminosity-to-gas mass ratios which are observed. These two facts make it possible to create a remnant with a higher central phase-space density than the precursors, removing one of the objections to creating ellipticals from mergers of spirals.

As discussed by Scoville & Soifer (1991), the ratio of space densities at the present epoch of infrared interacting galaxies and optical spiral galaxies is ∼ 2 × 10^{−3}. It is reasonable to believe that the duration of the infrared/starburst phase is approximately equal to the dynamical time (∼ 2 × 10^{8} years). Since two galaxies must take part in the interaction, the present day frequency of interaction-induced starbursts must be such that approximatley 2% of the present spiral galaxies would have passed through this phase in the last 10^{9} years. And since the IR luminosity to gas mass ratio (see §3) is enhanced by an order of magnitude during the starburst we conclude that a significant fraction (10-20%) of the stars now seen in spiral galaxies could have formed during interaction-induced starbursts. Including the effects of cosmological evolution on galaxy interaction rates makes this conclusion that much stronger.

A second major effect of galactic interactions is the redistribution and mixing of the ISM at different radii. Thus, ultimately the metalicity gradients in galaxies may be limited by galactic encounters such that more uniform metalicities will be observed in galaxies which have undergone relatively recent interactions. In addition, although the initial effect of the interaction on the low metalicity outer disk gas is to remove this gas into extended tidal features, much of this gas may eventually be absorbed back into the disk after several orbited periods. The result will be the gradual replenishment of the inner disk with low metalicity outer disk gas. This replenishment of the inner galactic disk both during and long after the galactic interaction may ultimately lead to a highly dissipative, “cold” disk out of which the next generation of post-interaction disk population stars will form. Clearly, the usual argument that interaction can not have played a significant role in the recent evolution of disk galaxies since the interaction would produce a vertically extended, hot disk population is no longer a strong argument if the majority of the thin disk stellar population was formed after the last interaction out of the resulting thin, gaseous disk.

Acknowledgement

This research was supported in part by NSF grants AST 90-16404 (NS and MSY) and AST 89-17744 and AST 90-23254 (JEH and JHvG). It is a pleasure to thank Alicia Rodriguez for the preparation of this manuscript.

REFERENCES

Alloin, D., Barvainis, R., Gordon, M.S. & Antonucci, R.R.J. 1992, Astron. Astro-

phys., 265, 429. Antonucci, R.R.J. & Miller, J.S. 1985, Ap.J., 297, 621. Barnes, J.E. 1988, Ap.J., 331, 699. Barnes, J.E. & Hernquist, L. 1991, Ap.J (Letters), 370, L65. Barvainis, R., Alloin, D. & Antonucci, R. 1989, Ap.J. (Letters), 337, L69. Boroson, T.A. & Oke, J.B. 1984, Ap.J., 337, 535. Brown, R.L. & Vanden Bout, P.A. 1991, A.J., 102, 1956. Bryant, P. & Scoville, N.Z. 1993, (in preparation). Carico, D.P., Graham, J.R., Matthews, K., Wilson, T.D., Soifer, B.T., Neugebauer,

G. & Sanders, D.B. 1990, Ap.J. (Letters), 349, L39. Carlstrom, J. 1990, unpublished PhD Thesis, Univesity of California (Berkeley). Condon, J.J., Condon, M.A., Gisler, G., & Puschell, J.J. 1982, Ap.J., 252,102. Dahari, O. 1985, Ap.J. Suppl., 57, 643. Hernquist, L. 1989 Nature, 340, 687. Hibbard, J.E. & van Gorkom, J.H. 1993, ASP conference series, Vol. 48, “The

Globular Cluster — Galaxy Connection”, eds. G.H. Smith & J.P. Brodie, 619. Hibbard, J.E., van Gorkom, J.H., Kasow, S., Westpfahl, D.J. 1993a, proceedings of the

third Tetons Summer school, “The Evolution of Galaxies and their Environment”, eds. J.M. Shull & H.A. Thronson (Dordrecht: Kluwer Academic Publishers) in press.

Hibbard, J.E., Guhathakurta, P., van Gorkom, J.H., Schweizer, F. 1993b, A.J. sub-

mitted. Hummel, E. & van der Hulst, J.M. 1986, Astr. Ap., 155, 151. Keel et al. 1985, Astr.J., 90, 708. Kikumoto, T., Tanigushi, Y., Suzuki, M., & Tomisada, K. 1992, “Millimeter and

Submillimeter Interferometry”, ASP Conference Series, (in press). Kormendy, J. & Sanders, D.B 1992, Ap.J. (Letters), 390, L53. Lawrence, A., Walker, D., Rowan-Robinson, M., Leech, K.J. & Penston, M.V. 1986,

M.N.R.A.S., 219, 687. Lo, K.Y., Cheung, D.W., Masson, C.R., Phillips, T.G., Scott, S.L. & Woody, D.P.

1987, Ap.J., 312, 574.

Matthews, K., Soifer, B.T., Graham, J.R. & Neugebauer, G. 1993, in preparation. Matthews, T.A. & Sandage, A.R. 1963, Ap.J., 138, 30. Nakai, N. Hayashi, M., Handa, T. Sofue, Y., Hasegawa, T. 1986, P.A.S.J., 38, 603. Neugebauer, G., Soifer, B.T., & Miley, G.R. 1985, Ap.J. (Letters), 295, L27. Norman, C.A. & Scoville, N.Z. 1988, Ap.J., 332, 124. Norris, R.P. 1988, M.N.R.A.S., 230, 345. Norris, R.P. Allen, D.A., Sramek, R.A., Kesteven, M.J., & Troup, E.R. 1990, Ap.J.,

359, 291. Okumura, S.K., Kawabe, R., Ishiguro, M., Kasuga, T., Morita, K.I. & Ishizuki, S.

1991, IAU Symposium No. 146, “Dynamics of Galaxies and their Molecular Distributions”, ed. F. Combes (Dordrecht: Kluwer), 425.

Peebles, P.J.E. 1980, The Large-Scale Structure of the Universe, (Princeton University

Press). Planesas, P. Mirabel, I.F. & Sanders, D.B. 1991 Ap.J., 370, 172. Planesas, P., Scoville, N.Z. & Myers, S.T. 1991, Ap.J., 369, 364. Rubin, V.C., Ford, W.K. & D’Odorico, S. 1970, Ap.J., 160, 801. Sanders, D.B., Scoville, N.Z. & Soifer, B.T. 1988, Ap.J. (Letters), 335, L1. Sanders, D.B., Scoville, N.Z. & Soifer, B.T. 1991, Ap.J., 370, 158. Sanders, D.B., Scoville, N.Z., Sargent, A.I. & Soifer, B.T. 1988a, Ap.J. (Letters),

324, L55. Sanders, D.B., Soifer, B.T., Elias, J.H., Madore, B.F., Matthews, K., Neugebauer,

G. & Scoville, N.Z. 1988b, Ap.J., 325, 74. Sargent, A.I. & Scoville, N.Z. 1991, Ap.J. (Letters), 366, L1. Schweizer, F. 1982, Ap.J., 252, 455. Schweizer, F. 1990, “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Spring-

er, Heidelberg), 60. Scoville, N.Z. & Good, J.C. 1989, Ap.J., 339, 149. Scoville, N.Z., Sanders, D.B., Sargent, A.I., Soifer, B.T. & Tinney, C.G. 1989, Ap.J.

(Letters), 345, L25. Scoville, N.Z., Sargent, A.I., Sanders, D.B. & Soifer, B.T. 1991, Ap.J. (Letters), 366,

L5. Scoville, N.Z., & Soifer, B.T. 1991, in “Massive Stars and Starbursts”, Space Telescope

Science Institute Symposium Series No. 5, ed. C. Leitherer, N.R. Wallborn, T.M. Heckman, and C.A. Norman (Cambridge University Press, Cambridge), 233.

Soifer, B.T., Sanders, D.B., Madore, B.F., Neugebauer, G., Danielson, G.E., Elias,

J.H., Lonsdale, C.J. & Rice, W.L. 1987, Ap.J., 320, 238. Solomon, P.M., Radford, S.J.E. & Downes, D. 1990, Ap.J. (Letters), 348, L53. Stanford, S.A., Balcells, M. 1991, Ap.J., 370, 118. Stanford, S.A., Bushouse, H.A. 1991, Ap.J., 371, 92. Stanford, S.A., Sargent, A.I., Sanders, D.B. & Scoville, N.Z. 1990 Ap.J., 349, 492. Stockton, A. & Ridgway, S.E. 1991, A.J., 102, 488. Suchkov, A., Allen, R.J., & Heckman, T.M. 1993, (preprint).

Tilanus, R.P. & Veilleux, S. 1993, (in preparation). Tinney, C.G., Scoville, N.Z., Sanders, D.B., & Soifer, B.T. 1990, Ap.J., 362, 473. Toomre, A. 1977, “The Evolution of Galaxies and Stellar Populations”, eds. B.M.

Tinsley & R.B. Larson (Yale Univ. Observatory), 401. Toomre, A. & Toomre, J. 1972, Ap.J, 178, 623. van Gorkom, J.H. 1993,proceedings of the third Tetons Summer school, “The Evo-

lution of Galaxies and their Environment”, eds. J.M. Shull & H.A. Thronson (Dordrecht: Kluwer Academic Publishers) in press.

Wang, Z., Scoville, N.Z., & Sanders, D.B. 1991, Ap.J., 368, 112. Wang, Z., Schweizer, F.P. & Scoville, N.Z. 1992, Ap.J., 396, 510. Wevers, B.H.M.R., van der Kruit, P.C. & Allen, R.J. 1986, Astro. Astrophys. Suppl.,

66, 505. White, S.D.M., & Rees, M. 1978, M.N.R.A.S., 183, 341. Wright, G.S., Joseph, R.D., Robertson, N.A., James, P.A. & Meikle, W.P.S. 1988,

M.N.R.A.S., 233, 1. Young, J.S., Claussen, M.J. & Scoville, N.Z. 1988, Ap.J., 324, 115. Young, J.S., Claussen, M.J., Kleinman, S.G., Rubin, V.C., & Scoville, N.Z. 1988,

Ap.J. (Letters), 331, L81. Young, J.S. & Scoville, N.Z. 1984, Ap.J., 287, 153. Yun, M.S., Ho, P.T.P. & Lo, K.Y. 1993a, Nature, (submitted). Yun, M.S., Ho, P.T.P. & Lo, K.Y. 1993b, Ap.J. (Letters), (in press). Yun, M.S. 1992, unpublished PhD Thesis (Harvard University). Yun, M.S. & Scoville, N.Z. 1993, (in preparation). Zwicky, F. 1956, Ergebnisse der Exakten Naturwissenschaften29, 344.